07:13
  • Monitoring and snoozing.
  • Still hanging out.
07:40
  • Alfio is unifying the SX and the DX adsec code. Right now, they have an error on DX, "Not Live"
  • They fixed that problem, now they are trying to fix the pupil problems on Bin2
  • /home/telescope/display_reset.sh
11:46
  • Lots of debugging happened. They got the bin2 pupils to work and now they are working on the pupil tracking routine, so that it works better for faint targets.
-- GregoryTaylor - 19 Feb 2019
Topic revision: r2 - 19 Feb 2019, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback